OpenStudy (anonymous):

Find the domain. Use interval notation. h(t) = sq rt of 4-3t

OpenStudy (akashdeepdeb):

Okay so domain in this case are the t values that would be applicable to find a value for h(t) So we have [/sqrt(4-3t)/] And we know that anything under square root should always be greater than or equal to 0 otherwise it will become a complex number!! So we can write the restrictions as 4-3t>=0 So To solve these kinds of inequalities to directly find the answer we ALWAYS try to write the co-efficient of the variable as 1 and we factorize the expression! So it can be also written as... 3(4/3-T)>=0 Or t<=4/3 So now T has to be lesser than or equal to 4/3 So the Domain HAS TO be (-infinity,4/3] (OBSERVE THE open and closed brackets] understood the concept behind this? :)

OpenStudy (anonymous):



OpenStudy (mathstudent55):

\(4 - 3t \ge 0\) Subtract 4 from both sides: \( - 3t \ge -4\) Divide both sides by -3. Remember to flip the inequality sign since you are dividing by a negative number: \( \dfrac{- 3t}{-3} \le \dfrac{-4}{-3}\) \( t \le \dfrac{4}{3}\) D: \( (-\infty, \frac{4}{3}] \)